# A New Algorithm for Whitney Stratification of Varieties

### v1

@@ Abstract.
We describe a new algorithm to compute Whitney stratifications of real and complex algebraic varieties. This algorithm is a modification of the algorithm of Helmer and Nanda (HN) [7, 9], but is made more efficient by using techniques for equidimensional decomposition rather than computing the set of associated primes of a polynomial ideal at a key step in the HN algorithm. We note that this modified algorithm may fail to produce a minimal Whitney stratification even when the HN algorithm would produce a minimal stratification. We, additionally, present an algorithm to coarsen any Whitney stratification of a complex variety to a minimal Whitney stratification; the theoretical basis for our approach is a classical result of Teissier.
@@ 1. Introduction
First introduced by Whitney in [15], the concept of Whitney stratification is now an essential tool in the study of singular spaces. This construction provides the basic structure needed to decompose singular spaces into smooth manifolds which join together in a desirable way. In this note we present a new method to compute such statifications which substantially improves on the state of the art in many cases, building on the recent works of [7, 9] by incorporating ideas used in symbolic algorithms which compute equidimensional decompositons of polynomial ideals, e.g. [2].
@@
in variables $x,y$ with parameter $z$. For $z<1$ the real variety $f_{z}=0$ is a (connected) nodal cubic, for $z>1$ the real variety $f_{z}=0$ has two connected components and is smooth, and at $z=0$ the curve is a cubic cusp, see Figure 1.
@@
To understand the geometry underlying this change in topology at $z=1$ for the parametric curve $f_{z}$ consider the variety $X=\mathbf{V}(f)$ in $\mathbb{R}^{3}$ defined by the same polynomial $f(x,y,z)=(y-1)^{2}-(x-z)(x-1)^{2}$, plotted in Figure 2. The surface $X$ is singular along the entire line ${\rm Sing}(X)=\mathbf{V}(x-1,y-1)$ and the projection map $\pi:\mathbb{R}^{3}\to\mathbb{R}$ onto the last coordinate, $(x,y,z)\to z$, restricted to either the manifold $X-\mathbf{V}(x-1,y-1)$ or to the manifold $\mathbf{V}(x-1,y-1)$ is a submersion, however the behavior of the fibers of $\pi$ restricted to $X-\mathbf{V}(x-1,y-1)$ depends on whether $z$ is greater than or less than $1$, while the behavior of the fibers of the restriction to $\mathbf{V}(x-1,y-1)$ is independent of $z$.
Figure 2. The surface $X=\mathbf{V}\left((y-1)^{2}-(x-z)(x-1)^{2}\right)$ in $\mathbb{R}^{3}$. It’s singular locus is $Y=\mathbf{V}(x-1,y-1)$ and its Whitney stratification arises from the flag $\{(1,1,1)\}\subset Y\subset X$. The dimension 2 stratum is $X-Y$, the two dimension 1 strata are the connected components of $Y-\{(1,1,1)\}$, and the dimension 0 stratum is $(1,1,1)$.
@@
More precisely a classical result known as Thom’s Isotopy Lemma [11, Proposition 11.1] allows us to partition the codomain of (proper) maps into regions of constant topology; Whitney’s condition is a critical component of this construction as illustrated by the example above  [^1] Note that in fact the map $\pi$ in the example is not proper, so Thom’s Isopotpy lemma does not hold for the affine map, however we can see that the result still holds for this example by passing to the projective closure in $\mathbb{P}^{2}\times\mathbb{C}$, where the corresponding map is proper..
@@
[…]
@@
This paper is organized as follows: in Section 2 we briefly review the necessary background on Gröbner bases, Whitney stratifications and local polar varieties. The main results, and the resulting algorithm for computing the Whitney stratification of a variety and for minimizing a given Whitney stratification, are presented in Section 3. Finally in Section 4 we give run time comparisons.
@@ 2.1. Gröbner Bases
A maximally independent subset $\mathbf{u}\subset\mathbf{x}$ of an ideal $I\subset\mathbb{K}[\mathbf{x}]$ is a set such that the map $\mathbb{K}[\mathbf{y}]\rightarrow\mathbb{K}[\mathbf{x},\mathbf{y}]/I$ is injective and such that the cardinality of $\mathbf{u}$ is equal to $\dim(I)$.
@@ 2.2. Whitney Stratifications
Throughout the paper, let $\mathbb{C}^{n}$ be the $n$-dimensional affine space over the complex numbers $\mathbb{C}$, $\mathbb{P}^{n}$ be the $n$-dimensional projective space over $\mathbb{C}$ with dual projective space $(\mathbb{P}^{n})^{*}$, and let $\mathbf{Gr}(k,n)$ denote the projective variety of $k$-dimensional vector spaces in $\mathbb{C}^{n}$. We assume throughout the rest of the paper that all varieties in $\mathbb{C}^{n}$ are defined over $\mathbb{Q}$, i.e. cut out by $n$-variate polynomials with coefficients in $\mathbb{Q}$. For any variety $X\subset\mathbb{C}^{n}$ we will denote by $I(X)$ the radical ideal of all polynomials vanishing on $X$, whereas $I_{X}$ will denote any ideal with $\mathbf{V}(I_{X})=X$, and in particular if $X$ is non-reduced (i.e. a scheme) $I_{X}$ will denote the ideal defining the scheme $X$.
@@
Let $Y\subset X$ be algebraic varieties in $\mathbb{C}^{n}$ with $\dim(Y)<\dim(X)$ and with $X-Y$ smooth. Let $Y_{\rm reg}$ denote the smooth part of $Y$. The pair $(X,Y)$ satisfies Whitney’s condition (B) if for any $p\in Y_{\rm reg}$
@@ A Whitney stratification of $X$ is a flag
In [7] an algebraic criterion based on primary decomposition of polynomial ideals is given that allows to check computationally if a given pair $X,Y\subset\mathbb{C}^{n}$ satisfies condition (B), and as such, to compute a Whitney stratification of a given variety by succesively computing singular loci and then checking condition (B) for all pairs of resuling varieties. The central contribution of this present work is to still apply the same algebraic criterion but to give a simple way in which having to explicitly compute primary decompositions can nonetheless be avoided, see Section 3.1.
@@
$\{(p,\zeta)\;|\;p\in X_{\text{reg}}\text{ and }T_{p}X_{\text{reg}}\subset\zeta\}.$
@@
This notion was used in [7] to give the aforementioned algebraic criterion for condition (B), which we restate below.
@@ 2.3. Local Polar Varieties and Minimal Whitney Stratifications
Let $X\subset\mathbb{C}^{n}$ be a variety of dimension $d$ and consider the conormal space $\mathbf{Con}(X)\subset\mathbb{C}^{n}\times\mathbb{P}^{n-1}$ with associated conormal map $\kappa_{X}:\mathbf{Con}(X)\to X$. For fixed $i$ with $11\leq i\leq d-$ 1, consider a dimension $d+i-1$ linear space in $\mathbb{C}^{n}$ with dual $L_{i}\subset(\mathbb{P}^{n-1})^{*}$. We use the convention $L_{0}=\mathbb{C}^{n}\times(\mathbb{P}^{n-1})^{*}$.
@@
Making this theorem computationally profitable will require us to compute sequences $m_{\bullet}(X,y)$ where $y$ lies in a subvariety $Y\subset X$. We will show how to do this in Section 3.3.
3. The New Whitney Stratification Algorithm
In this section we present the main results of this work, namely a new algebraic condition to identify subvarieties where condition (B) fails, new algorithms to compute a Whitney stratification of an algebraic variety and an algorithm to produce the unique minimal Whitney stratification for complex varieties along with theoretical results which guarantee the correctness of our algorithms. Our new formulation of Condition B is given in Section 3.1, the resulting stratification is given in Section 3.2, and finally an algorithm to coarsen a given Whitney stratification to the unique minimal stratification is given in Section 3.3.
@@ 3.1. A New Algebraic Criterion for Checking Whitney’s Condition
Throughout this section we denote by $\mathbb{C}[\mathbf{x}]$ the underlying polynomial ring of $\mathbb{C}^{n}$. We will study the problem of effectively computing a Whitney stratification of a complex affine algebraic variety $X\subset\mathbb{C}^{n}$. We first show how to, in the setting of Theorem 2.11, compute a superset of $A$ without having to compute a primary decomposition of $I_{Y}+I(\mathbf{Con}(X))$. For this we denote by $\mathbb{C}[\mathbf{x},\mathbf{y}]$ the underlying polynomial ring of $\mathbb{C}^{n}\times\mathbb{P}^{n-1}$, where $\mathbb{P}^{n-1}$ denotes the $(n-1)$-dimensional projective space over $\mathbb{C}$.
@@
Let $P\subset\mathbb{C}[\mathbf{x}]$ be a prime ideal and let $\mathbf{u}\subset\mathbf{x}$ be of cardinality $d$. If $P\cap\mathbb{C}[\mathbf{u}]=0$, then $\dim(P)\geq d$.
@@
Using Theorem 2.11 we obtain an alternative algebraic criterion which will also allow us to construct a Whitney stratification. When reading the result below keep in mind that $\mathbf{x},\mathbf{u},$ and $\mathbf{y}$ (and hence $\mathbf{x}-\mathbf{u}$) denote sets of variables.
@@
Further, clearly, $(J:h^{\infty})\cap\mathbb{C}[\mathbf{x}]=(I_{Y}:h^{\infty})$. For $P$ a minimal prime over $(I_{Y}:h^{\infty})$, choose a minimal prime $Q$ over $(J:h^{\infty})$ with $Q\cap\mathbb{C}[\mathbf{x}]=P$. Then we have
@@
and therefore $\dim(P)\geq\dim(Y)$ by Proposition 3.1. On the other hand, $P$ is also minimal over $I_{Y}$. Therefore $\dim(P)\leq\dim(Y)$ and finally $\dim(P)=\dim(Y)$. This shows that $Y-\mathbf{V}(h)$ is equidimensional of dimension equal to $\dim(Y)$.
@@
then we have $A\subset\mathbf{V}(h)$. By Theorem 2.11, the pair $(X,Y-(A\cup Y_{\text{sing}}))$ satisfies Whitney’s condition (B) and since $Y-\mathbf{V}(h)\subset Y-A$, so does $(X,Y-(\mathbf{V}(h)\cup Y_{\text{sing}}))$. ∎
3.2. The Algorithm
@@
Now we can give the following algorithm subroutine WhitPoints, which given any pair of varieties $X$ and $Y$ with $Y\subset X_{\text{sing}}$, returns a polynomial $h$ as in Theorem 3.2. This subroutine is the centerpiece of our algorithm for computing Whitney stratifications. As previously mentioned, we avoid the computation of associated primes in the conormal space of $X$ which an immediate application of Theorem 2.11 would require.
WhitPoints$(X,Y)$
@@
Note that, thanks to Proposition 2.5, a Gröbner basis of an ideal $I\subset\mathbb{C}[\mathbf{x}]$ with respect to a monomial order eliminating $\mathbf{x}-\mathbf{u}$, where $\mathbf{u}\subset\mathbf{x}$, is also a Gröbner basis of $I\mathbb{C}(\mathbf{u})[\mathbf{x}-\mathbf{u}]$. The stated result then follows immediately from Theorem 3.2. ∎
@@
Note that, as Theorem 3.2 does not rely on $Y$ being irreducible, it is not strictly needed for the correctness of our algorithm to represent the occuring varieties by their $\mathbb{Q}$-irreducible components. We chose to present the algorithm like this here for ease of reading. It would not be difficult to adapt Algorithm 5 in [2] to our situation, then we would have to perform only equidimensional decomposition instead of finding the minimal primes over a given ideal which is generally much easier.

### v2

# Faster Computation of Whitney Stratifications and Their Minimization

@@ Abstract.
We describe two new algorithms for the computation of Whitney stratifications of real and complex algebraic varieties. The first algorithm is a modification of the algorithm of Helmer and Nanda (HN) [7, 9], but is made more efficient by using techniques for equidimensional decomposition rather than computing the set of associated primes of a polynomial ideal at a key step in the HN algorithm. We note that this modified algorithm may fail to produce a minimal Whitney stratification even when the HN algorithm would produce a minimal stratification.
The second algorithm coarsens a given Whitney stratification of a complex variety to the unique minimal Whitney stratification; we refer to this as the minimization of a stratification. The theoretical basis for our approach is a classical result of Teissier. To our knowledge this yields the first algorithm for computing a minimal Whitney stratification.
Keywords. Whitney stratification, Gröbner basis, Conormal variety.
The second author is the corresponding author.
@@ 1. Introduction
First introduced by Whitney in [20], the concept of Whitney stratification is now an essential tool in the study of singular spaces. This construction provides the basic structure needed to decompose singular spaces into smooth manifolds which join together in a desirable way. In this note we present a new method to compute such statifications which substantially improves on the state of the art in many cases, building on the recent works of [7, 9] by incorporating ideas used in symbolic algorithms which compute equidimensional decompositons of polynomial ideals, e.g. [2]. We additionally present an algorithm for minimizing a given Whitney stratification.
@@
in variables $x,y$ with parameter $z$. For $z<1$ the real variety $f_{z}=0$ is a (connected) nodal cubic, for $z>1$ the real variety $f_{z}=0$ has two connected components and is smooth, and at $z=0$ the curve is a cubic cusp, see Figure˜1.
@@
To understand the geometry underlying this change in topology at $z=1$ for the parametric curve $f_{z}$ consider the variety $X=\mathbf{V}(f)$ in $\mathbb{R}^{3}$ defined by the same polynomial $f(x,y,z)=(y-1)^{2}-(x-z)(x-1)^{2}$, plotted in Figure˜2. The surface $X$ is singular along the entire line ${\rm Sing}(X)=\mathbf{V}(x-1,y-1)$ and the projection map $\pi:\mathbb{R}^{3}\to\mathbb{R}$ onto the last coordinate, $(x,y,z)\to z$, restricted to either the manifold $X-\mathbf{V}(x-1,y-1)$ or to the manifold $\mathbf{V}(x-1,y-1)$ is a submersion, however the behavior of the fibers of $\pi$ restricted to $X-\mathbf{V}(x-1,y-1)$ depends on whether $z$ is greater than or less than $1$, while the behavior of the fibers of the restriction to $\mathbf{V}(x-1,y-1)$ is independent of $z$.
Figure 2. The surface $X=\mathbf{V}\left((y-1)^{2}-(x-z)(x-1)^{2}\right)$ in $\mathbb{R}^{3}$. It’s singular locus is $Y=\mathbf{V}(x-1,y-1)$ and its Whitney stratification arises from the flag $\{(1,1,1)\}\subset Y\subset X$. The dimension 2 stratum is XYX-Y, the two dimension 1 strata are the connected components of $Y-\{(1,1,1)\}$, and the dimension 0 stratum is $(1,1,1)$.
@@
More precisely a classical result known as Thom’s Isotopy Lemma [12, Proposition 11.1] allows us to partition the codomain of (proper) maps into regions of constant topology; Whitney’s condition is a critical component of this construction as illustrated by the example above  [^1] Note that in fact the map $\pi$ in the example is not proper, so Thom’s Isopotpy lemma does not hold for the affine map, however we can see that the result still holds for this example by passing to the projective closure in $\mathbb{P}^{2}\times\mathbb{C}$, where the corresponding map is proper..
@@
$\mathbf{V}\left(x_{4},\,x_{2},\,x_{1}\right),\>\mathbf{V}\left(x_{2},\,x_{1},\,4\,x_{3}^{3}-27\,x_{4}^{2}\right),\>\mathbf{V}\left(x_{2},\,x_{1}-x_{3},\,x_{3}^{2}-x_{4}\right).$ (2)
@@
This paper is organized as follows: in Section˜2 we briefly review the necessary background on Gröbner bases, Whitney stratifications and local polar varieties. The main results, and the resulting algorithm for computing the Whitney stratification of a variety and for minimizing a given Whitney stratification, are presented in Section˜3. Finally in Section˜4 we give run time comparisons.
Acknowledgements
[…]
@@ 2.1. Gröbner Bases
A maximally independent subset $\mathbf{u}\subset\mathbf{x}$ of an ideal $I\subset\mathbb{K}[\mathbf{x}]$ is a set such that the map $\mathbb{K}[\mathbf{u}]\rightarrow\mathbb{K}[\mathbf{x}]/I$ is injective and such that the cardinality of $\mathbf{u}$ is equal to $\dim(\mathbb{K}[\mathbf{x}]/I)$.
@@ 2.2. Whitney Stratifications
Throughout the paper, let $\mathbb{C}^{n}$ be the $n$-dimensional affine space over the complex numbers $\mathbb{C}$, $\mathbb{P}^{n}$ be the $n$-dimensional projective space over $\mathbb{C}$ with dual projective space $(\mathbb{P}^{n})^{*}$, and let $\mathbf{Gr}(k,n)$ denote the projective variety of $k$-dimensional vector spaces in $\mathbb{C}^{n}$. We assume throughout the rest of the paper that all varieties in $\mathbb{C}^{n}$ are defined over $\mathbb{Q}$, i.e. cut out by $n$-variate polynomials with coefficients in $\mathbb{Q}$. For any variety $X\subset\mathbb{C}^{n}$ we will denote by $I(X)$ the radical ideal of all polynomials vanishing on $X$. We will frequently use the notation $I_{X}$ to denote any ideal with $\mathbf{V}(I_{X})=X$. Note that such an ideal $I_{X}$ is of course not uniquely defined.
@@
Let $Y\subset X$ be algebraic varieties in $\mathbb{C}^{n}$ with $\dim(Y)<\dim(X)$ and with XYX-Y smooth. Let $Y_{\rm reg}$ denote the smooth part of $Y$. The pair $(X,Y)$ satisfies Whitney’s condition (B) if for any $p\in Y_{\rm reg}$
@@ A Whitney stratification of $X$ is a flag
In [7] an algebraic criterion based on primary decomposition of polynomial ideals is given that allows to check computationally if a given pair $X,Y\subset\mathbb{C}^{n}$ satisfies condition (B), and as such, to compute a Whitney stratification of a given variety by successively computing singular loci and then checking condition (B) for all pairs of resulting varieties. The central contribution of this present work is to still apply the same algebraic criterion but to give a simple way in which having to explicitly compute primary decompositions can nonetheless be avoided, see Section˜3.1.
@@
$\{(p,\zeta)\;|\;p\in X_{\text{reg}}\text{ and }T_{p}X_{\text{reg}}\subset\zeta^{\perp}\}.$
@@
This notion was used in [7] to give the aforementioned algebraic criterion for condition (B), which we restate below [^2] We note that in [7] this theorem is stated and proved (in the correction [8] to [7]) for the case where $X$ (and hence $Y$) are projective varieties. However the restriction to the projective case is not necessary and the same theorem holds, with exactly the same proof in the affine case (specifically the proof of Theorem 2.1 in [8], which is the corrected version of Theorem 4.3 in [7]) . Hence we do not redo the proof here and simply re-phrase the final result to the affine setting..
@@ 2.3. Local Polar Varieties and Minimal Whitney Stratifications
Let $X\subset\mathbb{C}^{n}$ be a variety of dimension $d$ and consider the conormal space $\mathbf{Con}(X)\subset\mathbb{C}^{n}\times\mathbb{P}^{n-1}$ with associated conormal map $\kappa_{X}:\mathbf{Con}(X)\to X$. For fixed $i$ with $1\leq i\leq d-1$, consider a dimension d+i1d+i-1 linear space in $\mathbb{C}^{n}$ with dual $L_{i}\subset(\mathbb{P}^{n-1})^{*}$. We use the convention $L_{0}=\mathbb{C}^{n}\times(\mathbb{P}^{n-1})^{*}$.
@@
Making this theorem computationally profitable will require us to compute sequences $m_{\bullet}(X,y)$ where $y$ lies in a subvariety $Y\subset X$. We will show how to do this in Section˜3.3.
3. New Whitney Stratification Algorithms
In this section we present the main results of this work, namely a new algebraic condition to identify subvarieties where condition (B) fails, new algorithms to compute a Whitney stratification of an algebraic variety and an algorithm to produce the unique minimal Whitney stratification for complex varieties along with theoretical results which guarantee the correctness of our algorithms. Our new formulation of Condition B is given in Section˜3.1, the resulting stratification is given in Section˜3.2, and finally an algorithm to coarsen a given Whitney stratification to the unique minimal stratification is given in Section˜3.3.
@@ 3.1. A New Algebraic Criterion for Checking Whitney’s Condition
Throughout this section we denote by $\mathbb{C}[\mathbf{x}]$ the underlying polynomial ring of $\mathbb{C}^{n}$. We will study the problem of effectively computing a Whitney stratification of a complex affine algebraic variety $X\subset\mathbb{C}^{n}$. We first show how to, in the setting of Theorem 2.11, compute a superset of $A$ without having to compute a primary decomposition of $I_{Y}+I(\mathbf{Con}(X))$. For this we denote by $\mathbb{C}[\mathbf{x},\mathbf{y}]$ the underlying polynomial ring of $\mathbb{C}^{n}\times\mathbb{P}^{n-1}$, where $\mathbb{P}^{n-1}$ denotes the (n1)(n-1)-dimensional projective space over $\mathbb{C}$.
@@
Let $P\subset\mathbb{C}[\mathbf{x}]$ be a prime ideal and let $\mathbf{u}\subset\mathbf{x}$ be of cardinality $d$. If $P\cap\mathbb{C}[\mathbf{u}]=0$, then $\dim(\mathbb{C}[x]/P)\geq d$.
@@
Using Theorem˜2.11 we obtain an alternative algebraic criterion which will also allow us to construct a Whitney stratification. When reading the result below keep in mind that $\mathbf{x},\mathbf{u},$ and $\mathbf{y}$ (and hence $\mathbf{x}-\mathbf{u}$) denote sets of variables.
@@
Note that $\sqrt{J}\cap\mathbb{C}[\mathbf{x}]=I(Y)$ and hence, since $h\in\mathbb{C}[\mathbf{u}]$, $\sqrt{(J:h^{\infty})}\cap\mathbb{C}[\mathbf{x}]=\sqrt{(I_{Y}:h^{\infty})}$. Hence, for $P$ a minimal prime over $(I_{Y}:h^{\infty})$, there is a minimal prime $Q$ over $(J:h^{\infty})$ with $Q\cap\mathbb{C}[\mathbf{x}]=P$. Then we have
@@
and therefore $\dim(\mathbf{V}(P))\geq\dim(Y)$ by Proposition˜3.1. On the other hand, $P$ is also minimal over $I_{Y}$. Therefore $\dim(\mathbf{V}(P))\leq\dim(Y)$ and finally $\dim(\mathbf{V}(P))=\dim(Y)$. This shows that $Y-\mathbf{V}(h)$ is equidimensional of dimension equal to $\dim(Y)$.
@@
then we have $A\subset\mathbf{V}(h)$. By Theorem˜2.11, the pair $(X,Y-(A\cup Y_{\text{sing}}))$ satisfies Whitney’s condition (B) and since $Y-\mathbf{V}(h)\subset Y-A$, so does $(X,Y-(\mathbf{V}(h)\cup Y_{\text{sing}}))$. ∎
3.2. The Whitney Stratification Algorithm
@@
Now we can give the following algorithm subroutine WhitPoints, which given any pair of varieties $X$ and $Y$ with $Y\subset X_{\text{sing}}$, returns a polynomial $h$ as in Theorem˜3.2. This subroutine is the centerpiece of our algorithm for computing Whitney stratifications. As previously mentioned, we avoid the computation of associated primes in the conormal space of $X$ which an immediate application of Theorem˜2.11 would require.
WhitPoints(X,Y)(X,Y)
@@
Note that, thanks to Proposition 2.5, a Gröbner basis of an ideal $I\subset\mathbb{C}[\mathbf{x}]$ with respect to a monomial order eliminating $\mathbf{x}-\mathbf{u}$, where $\mathbf{u}\subset\mathbf{x}$, is also a Gröbner basis of $I\mathbb{C}(\mathbf{u})[\mathbf{x}-\mathbf{u}]$. The stated result then follows immediately from Theorem˜3.2. ∎
@@
Note that, as Theorem˜3.2 does not rely on $Y$ being irreducible, it is not strictly needed for the correctness of our algorithm to represent the occuring varieties by their $\mathbb{Q}$-irreducible components. We chose to present the algorithm like this here for ease of reading. It would not be difficult to adapt Algorithm 5 in [2] to our situation, then we would have to perform only equidimensional decomposition instead of finding the minimal primes over a given ideal which is generally much easier.